You are involved in the evaluation of viewership of television shows
for your local TV station. You have been tasked to establish a 95% confidence
interval for the proportion of viewers that recently watched an episode of a show called 'CSI Phoenix'.
You have randomly queried 400 local adults and determined that 150 of them watched the last episode of the show.
>> create a 95% confidence interval for the proportion of folks in the population that viewed the show (remember -- confidence intervals ... have lower and upper bounds)

Answer by stanbon(75887)

You have randomly queried 400 local adults and determined that 150 of them watched the last episode of the show.
>> create a 95% confidence interval for the proportion of folks in the population that viewed the show
--------------
Sample proportion = 150/400 = 15/40
---
standard error = E = 1.96*sqrt[(15/40)(25/40)/400] = 1.96*0.0242 = 0.0474
===============
95% CI: (15/40)-0.0474 < p < (15/40)+0.0474
--
95% CI: 0.3276 < p < 0.4224
===============================
Cheers,
Stan H.